OpenStudy (anonymous):

Solve the following equation. -4=g+1 A.g=-4 B.g=4 C.g=-5 D.g=-3 WILL GIVE MEDAL FOR CORRECT ANSWER!!!!

OpenStudy (austinl):

\(-4=g+1\) Subtract 1 from each side, \(g=-4-1=~?\)

OpenStudy (anonymous):

\[-4=g+1\]\[-5=g\] so the correct answer is C